OpenStudy (pythagoras123):

Find C and show your working (steps) \[\frac{C-100}{100}\] = \[\frac{2C-60}{C+60}\]

OpenStudy (inkyvoyd):

Multiply both sides by 100(c+60)

OpenStudy (inkyvoyd):

solve for C, should be a quadratic

OpenStudy (inkyvoyd):

C cannot equal -60, becaue that would mean dividing by 0

OpenStudy (saifoo.khan):

\[\frac{C-100}{100} =\frac{2C-60}{C+60}\]Cross multiply.

OpenStudy (saifoo.khan):

\[(C-100) (C+60) = (2C-60)(100)\]Solve.

OpenStudy (inkyvoyd):

Foo foo give better answer

OpenStudy (inkyvoyd):

But still, make sure C=/=-60

OpenStudy (saifoo.khan):

First let him solve @inkyvoyd

OpenStudy (inkyvoyd):

I mean, ifthe answer comes out as -60, it won't be accurate, because the denominator is never equal to zero.

OpenStudy (anonymous):

(c -100)/100 = (2c-60)/ (c+60) denominator are 100 and c+60 LCM = 100(c+60) multiply both sides with 100(c-60) (c-100)(c+60) = (2c-60)(100) (c-100)(c+60) =c^2 +60c -100c -6000 =c^2 - 40c -6000 (2c-60)(100) =200c -6000 (c-100)(c+60) = (2c-60)(100) c^2 - 40c -6000= 200c -6000 c^2 -40c - 6000 - 200c = 200c -200c-6000 c^2 -240c=0 c(c-240)=0 c=0 or c=240